# Analytical Solutions for Turbulent Channel Flow Using Alexeev and Navier-Stokes Hydrodynamic Equations: Comparison with Experiments

### v1

@@ Abstract
Keywords: Generalized Boltzmann Equation; Alexeev Hydrodynamic Equations; Navier-Stokes equations; Analytical solution; Numerical solution; Experimental validation; Turbulent channel flow; Turbulent boundary layer.
@@ 2 Generalized Boltzmann Equation
Alexeev assumed that the parameter $\tau$ defined in this way is the mean time between collisions [2](p.52). It was found that the parameter $\tau$ is a material property that, in the case of the hydrodynamic equations, gives rise to the length scale $\delta=\sqrt{\tau\nu}$, where $\nu$ is the kinematic viscosity. We will show that $\delta$ is a boundary layer thickness scale and that $\delta$ coincides with the Kolmogorov microscale observed in experiments. This $\delta$ has a value of approximately 0.6 mm for distilled water and approximately the same for air over a wide pressure range (to hundreds of atmospheres).
@@ 4.4 Validation for the turbulent solution $U^{T}$
Equation (37) results in a nonlinear system of finite-difference equations, which was solved by minimizing the squared residual using the Powell hybrid method (MINPACK) [29]. The same approach was used for to solve equation (38).
@@ 5.2 Comparison of solutions for pipe flow
In van Doorne and Westerweel (2007) experiment, a circular pipe with an inner diameter of 40 mm and a total length of 28 m was used, so the ratio of diameter to length was $D/L=700$ [7]. The working fluid was tap water. Due to a well designed contraction and thermal isolation of the pipe, the flow can be kept laminar up to Re = 60,000. It was found that the Coriolis force affects the velocity profile, as leads to an asymmetry for Re ¿ 3,000, as explained by Draad (1998) [8], so the orientation of the experimental pipe was chosen to exclude the Coriolis force. The measurements at Re=7200 were carried out at the distance of 26 m from the inlet, and stereoscopic-PIV was used [7]. The analytical solution (39) can be applied to circular pipe flow, when $\delta=\sqrt{\tau_{0}\nu}/L_{0}\ll 1$, which is the case, as $\delta=0.047$. Figure 5 shows the experimental data for the streamwise mean velocity digitized from [7] along with several plots: (i) Navier-Stokes solution, the laminar (parabolic) flow profile (green line), (ii) turbulent (superexponential) flow profile (blue line) and (iii) the AHE solution (red line). The value of $\gamma$ is $\gamma=0.68$, obtained from minimal viscous dissipation principle [12]. One can see that neither the laminar nor turbulent solution fit the data, but the superposition given by equation (39), the AHE solution (red line), provides good comparison to the experimental data.
@@
A series of 26 experiments was conducted using the Princeton Superpipe setup, [33] with air as the working fluid. The circular pipe diameter was 12.9 cm. The temperature was near ambient (295-300 K), and the pressure was varied between 1 and 186 atmospheres for 26 different experiments at Reynolds numbers, ranging from $3.15\cdot 10^{4}$ to $35.2\cdot 10^{6}$. The effects of the Coriolis force which potentially can affect the results, according to [7] and [8], were not discussed by the authors of [33].
@@ 6.2 Superposition coefficients $\gamma$
6.3 Boundary thickness scale $\delta$
The boundary thickness scale and the similarity parameter $\delta=\sqrt{\tau_{0}\nu}/L_{0}$ is the most important parameter for the turbulent flows, discovered in the analytical solution. The value of $\delta$ coincides with the Kolmogorov microscale, observed in experiment by Koseff and Street (1984) for 3D lid driven cavity flow [25].
@@ 7 Conclusion
The analytical solution of the Alexeev Hydrodynamic Equations for turbulent channel flow demonstrates significantly better agreement with the mean velocity experimental data than solutions obtained from the Navier-Stokes equations. These solutions successfully capture the correct velocity behavior across the entire turbulent boundary layer and into the external flow, spanning from the inner viscous sublayer to the outer layer of the boundary layer. The analytical solution reveals an important boundary layer thickness scale through the similarity parameter $\delta$, which coincides with the Kolmogorov microscale observed in experiments. Novel mechanisms for turbulence generation were identified and a turbulence control method was proposed.

### v2

@@ Abstract
Keywords: Generalized Boltzmann Equation; Alexeev Hydrodynamic Equations; Navier-Stokes equations; Turbulent channel flow; Turbulent boundary layer; Analytical solution; Numerical solution; Experimental validation.
@@ 2 Generalized Boltzmann Equation
Alexeev assumed that the parameter $\tau$ defined in this way is the mean time between collisions [2](p.52). This is not correct. It was found that the parameter $\tau$ is of much larger value, and is a material property, that, in the case of the hydrodynamic equations, gives rise to the length scale $\delta=\sqrt{\tau\nu}$, where $\nu$ is the kinematic viscosity. We will show that $\delta$ is a boundary layer thickness scale and that $\delta$ coincides with the Kolmogorov microscale observed in experiments. This $\delta$ has a value of approximately 0.6 mm for distilled water and approximately the same for air over a wide pressure range (to hundreds of atmospheres).
@@ 4.4 Validation for the turbulent solution $U^{T}$
Equation (37) results in a nonlinear system of finite-difference equations, which was solved by minimizing the squared residual using the Powell hybrid method (MINPACK) [29]. The same approach was used to solve equation (38).
@@ 5.2 Comparison of solutions for pipe flow
In van Doorne and Westerweel (2007) experiment, a circular pipe with an inner diameter of 40 mm and a total length of 28 m was used, so the ratio of diameter to length was $D/Length=700$ [7]. The working fluid was tap water. Due to a well designed contraction and thermal isolation of the pipe, the flow can be kept laminar up to Re = 60,000. It was found that the Coriolis force affects the velocity profile, as leads to an asymmetry for Re ¿ 3,000, as explained by Draad (1998) [8], so the orientation of the experimental pipe was chosen to exclude the Coriolis force. The measurements at Re=7200 were carried out at the distance of 26 m from the inlet, and stereoscopic-PIV was used [7]. The analytical solution (39) can be applied to circular pipe flow, when $\delta=\sqrt{\tau_{0}\nu}/L_{0}\ll 1$, which is the case, as $\delta=0.047$. Figure 5 shows the experimental data for the streamwise mean velocity digitized from [7] along with several plots: (i) Navier-Stokes solution, the laminar (parabolic) flow profile (green line), (ii) turbulent (superexponential) flow profile (blue line) and (iii) the AHE solution (red line). The value of $\gamma$ is $\gamma=0.68$, obtained from minimal viscous dissipation principle [12]. One can see that neither the laminar nor turbulent solution fit the data, but the superposition given by equation (39), the AHE solution (red line), provides good comparison to the experimental data.
@@
A series of 26 experiments was conducted using the Princeton Superpipe setup, with air as the working fluid [33]. The circular pipe diameter was 12.9 cm. The temperature was near ambient (295-300 K), and the pressure was varied between 1 and 186 atmospheres for 26 different experiments at Reynolds numbers, ranging from $3.15\cdot 10^{4}$ to $35.2\cdot 10^{6}$. The effects of the Coriolis force which potentially can affect the results, according to [7] and [8], were not discussed by the authors of [33].
@@ 6.2 Superposition coefficients $\gamma$
6.3 Boundary layer thickness scale $\delta$
The boundary layer thickness scale and the similarity parameter $\delta=\sqrt{\tau_{0}\nu}/L_{0}$ is the most important parameter for the turbulent flows, discovered in the analytical solution. The value of $\delta$ coincides with the Kolmogorov microscale, observed in experiment by Koseff and Street (1984) for 3D lid driven cavity flow [25].
@@ 7 Conclusion
The analytical solution of the Alexeev Hydrodynamic Equations for turbulent channel flow demonstrates significantly better agreement with the mean velocity experimental data than solutions obtained from the Navier-Stokes equations. These solutions successfully capture the correct velocity behavior across the entire turbulent boundary layer and into the external flow, spanning from the inner viscous sublayer to the outer layer of the boundary layer. The analytical solution reveals an important boundary layer thickness scale through the similarity parameter $\delta$, which coincides with the Kolmogorov microscale observed in experiments. The mechanisms for turbulence generation were identified and a turbulence control method was proposed.